OpenStudy (anonymous):

Slope equation: passes thru (7,1) parallel to y = -x + 3 I think for the equation it would be y = x (+/-) b? Not sure I can't reply on my phone but I will read over your replies Til I understand then ill medal you.

OpenStudy (anonymous):

Parallel lines have the same slope. The slope of the given line \(y=-x+3\) is \(-1\). Use the point-slope formula to find the equation of the line parallel to this one that contains (7,1): \[y-1=-1(x-7)\]